how can i find the vertex for y=2x2+7x+2
Answer by lwsshak3(11628)

how can i find the vertex for y=2x2+7x+2
Given equation is that of a parabola.
Its vertex form: y=A(x-h)^2+k, (h,k)=(x,y) coodinates of the vertex
complete the square:
y=2(x2+7x/2+49/16)-49/8+2
y=2(x+7/4)^2-33/8
vertex:(-7/4,-33/8)